# Faces of Fairness: Examining Bias in Facial Expression Recognition Datasets and Models

### v1

@@ 5 Discussion and Future Works
This research presented a comprehensive analysis of bias and fairness in FER. We examined four in-the-wild datasets (AffectNet, ExpW, Fer2013, and RAF-DB) and six popular deep models (MobileNet, ResNet, XceptionNet, ViT, CLIP, and GPT-4o-mini). We mainly focused on studying bias in demographic attributes, including age, gender, and race, as well as label distribution.
Based on the experiments conducted in this research, bias was examined in both datasets and models. Initially, bias within the datasets was analyzed using the Richness, Evenness, and Dominance metrics. The results indicated that ExpW is the most biased dataset, whereas AffectNet was identified as the least biased. Additionally, six state-of-the-art models were evaluated using four bias metrics: Equalized Odds, Equal Opportunity, Demographic Parity, and Treatment Equality. The findings revealed that, despite achieving high accuracy, the GPT-4o-mini and ViT models exhibited the highest levels of bias.
@@
This research highlighted the significance of considering bias and fairness in FER tasks. Establishing a foundation for enhancing the equity of FER models is essential. These improvements have the potential to benefit critical applications, including human-computer interaction (HCI), mental health monitoring, and surveillance. Future studies will focus on proposing methods to address bias in FER tasks while maintaining high accuracy.

### v2

@@ 5 Discussion and Future Works
Our analysis highlights that both the datasets and models used for facial expression recognition (FER) exhibit significant biases that could impact the fairness and generalizability of FER systems. Across the four in-the-wild datasets—AffectNet, Fer2013, RAF-DB, and ExpW—demographic imbalances are apparent. Race consistently emerges as the most biased attribute, while Age and Gender exhibit moderate but persistent disparities. Similarly, expression distributions are skewed, with Neutral and Happy overrepresented and Fear and Disgust underrepresented. These imbalances suggest that models trained on these datasets may inherit and even amplify existing biases.
The model evaluation further emphasizes the distinction between accuracy and fairness. While specialized FER models such as POSTER and CEPrompt achieve high overall accuracy, they do not effectively mitigate bias, particularly for underrepresented expressions and demographic groups. GPT-4o-mini and ViT, despite strong performance in some tasks, were identified as the most biased models, whereas residual-based architectures like ResNet and XceptionNet demonstrated relatively lower bias. Interestingly, some models partially compensated for imbalances in the data, reducing the impact of skewed distributions, but no model completely addressed the fairness challenges inherent to FER tasks.
These findings underline two important points. First, high accuracy alone is not sufficient to ensure equitable FER performance across all demographic groups and expressions. Second, addressing bias requires attention at both the dataset and model levels: collecting more balanced datasets, particularly for underrepresented expressions and demographic groups, is essential, as is developing models that explicitly account for these biases during training. Overall, our results call for a holistic approach that evaluates FER systems not only on predictive performance but also on fairness metrics, ensuring more reliable and equitable deployment in real-world applications.
@@
This research highlighted the significance of maintaining a balanced data distribution across demographic groups and ensuring the fairness of deep models. Establishing a foundation for enhancing the equity of FER models is essential. These improvements have the potential to benefit critical applications, including human-computer interaction (HCI), mental health monitoring, and surveillance. Future studies will focus on proposing methods to address bias in FER tasks while maintaining high accuracy.